OpenStudy (anonymous):

33.313 as a mixed number ?how?

OpenStudy (anonymous):

33.313 Take this number and multiply it by 1000 which is 33313. Put that number over 1000 since it basically equals the same number.

OpenStudy (anonymous):

\[\frac{ 33313 }{ 1000 }\]

OpenStudy (anonymous):

Questions?

OpenStudy (anonymous):

Can I have a question?@anisanox1

OpenStudy (anonymous):

sure?

OpenStudy (anonymous):

33 is an integer. 0.313 is in decimal form but can be expressed as a fraction in a form\[\frac{313}{1000}.\]that's why 33.313 is a mixed number. It is a mixture of integer and a fraction.

OpenStudy (anonymous):

\(\frac{ 33313 }{ 1000 } \)this guy is fraction, not mixed number. Am I right?

OpenStudy (anonymous):

ohhhh crap

OpenStudy (anonymous):

so it would look like 33\[33 \frac{ 313 }{ 1000 }\]

OpenStudy (anonymous):

yup you are right...

OpenStudy (anonymous):

Got you. Thanks

OpenStudy (anonymous):

hahaha whoops Ok continuing FROM the fraction that I inputted: Just divide 33313 by 1000 and you get 33 (Evenly without decimals) Now you take the rest of the number which is 313, and put that over 1000. \[33\frac{ 313 }{ 1000 }\]

OpenStudy (anonymous):

yay! okay um how would you do \[2\frac{ 3 }{ 10}\] as a decimal

OpenStudy (anonymous):

Well you take the 2 which is \[\frac{ 20 }{ 10 }\] Cuz they basically = the same thing. Then you add 3 from the fraction, \[\frac{ 23 }{ 10 }\] If you are dividing by 10 you are practically just taking the decimal point which is behind 23(.) and move it to the left once. \[2\frac{ 3 }{ 10 } = 2.3\]

OpenStudy (anonymous):

Similar with multiplying, move the decimal point to the right this time though.

OpenStudy (anonymous):

okay i understand more now. umm one more question sorry. im not good with fractions and other stuff with math. \[\frac{ 127 }{ 1000 }\] as a decimal number

OpenStudy (anonymous):

do i divide it

OpenStudy (anonymous):

Alright, as this goes on, I'm assuming that we can't use a calculator? 1000 has 3 zeroes and is in the denominator

OpenStudy (anonymous):

If it's in the denominator, it's basically telling you to divide the top number (numerator) with the bottom number (denominator)

OpenStudy (anonymous):

Dividing 3 zeroes would make your decimal point move 3 positions to the left \[0.0127\]

OpenStudy (anonymous):

THANK YOU!!!!! soooooooo much

OpenStudy (anonymous):

No problem :) Welcome to Open Study.

OpenStudy (anonymous):

er you're 33 so not exactly welcome...

OpenStudy (anonymous):

I am 18 years old

OpenStudy (anonymous):

No I meant the "SmartScore"

OpenStudy (anonymous):

Which turned into 35...

OpenStudy (anonymous):

Btw I'm 16, just to even out the playing field

OpenStudy (anonymous):

oh well i only asked a few questions but im not great with math or english or history so gotta get help somewhere

OpenStudy (anonymous):

haha alright.

hero (hero):

33.313 is read mathematically as thirty three and three hundred thirteen thousandths.

hero (hero):

Which as a fraction can immediately be written as \[33 + \frac{313}{1000}\]